40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Aug 15 14:10:56 2011, Jenne, Update, SUS, Monday SUS update 7x
    Reply  Mon Aug 15 17:36:10 2011, jamie, Update, SUS, Strangeness with ETMY (was: Monday SUS update) 
    Reply  Tue Aug 30 09:51:37 2016, Steve, Update, SUS, the chamber HEPA tents are back ETMYtentHEPA.jpgvented_for_60_days.png
       Reply  Wed Aug 31 23:01:02 2016, gautam, Update, SUS, the chamber HEPA tents are back 
          Reply  Thu Sep 1 15:40:52 2016, Steve, Update, SUS, HEPAs on each chamber door chamberDoorHepaEA.pdf
Message ID: 5241     Entry time: Mon Aug 15 17:36:10 2011     In reply to: 5239
Author: jamie 
Type: Update 
Category: SUS 
Subject: Strangeness with ETMY (was: Monday SUS update) 

For some reason ETMY has changed a lot.  Not only does it now have the worst "badness" (B matrix condition number) at ~10, but the frequency of all the modes have shifted, some considerably.  I did accidentally bump the optic when Jenne and I were adjusting the OSEMs last week, but I didn't think it was that much.  The only thing I can think of that would cause the modes to move so much is that the optic has been somehow reseated in it's suspension.  I really don't know how that would have happened, though.

Jenne and I went in to investigate ETMY, to see if we could see anything obviously wrong.  Everything looks to be ok.  The magnets are all well centered in the OSEMs, and the PDMon levels look ok.

We rechecked the balance of the table, and tweaked it a bit to make it more level.  We then tweaked the OSEMs again to put them back in the center of their range.  We also checked the response by using the lockin method to check the response to POS and SIDE drive in each of the OSEMs (we want large POS response and minimal SIDE response).  Everything looked ok.

We're going to take another freeswing measurement and see how things look now.  If there are any suggestions what should be done (if anything), about the shifted modes, please let us know.

ELOG V3.1.3-